# HybridNorm: Towards Stable and Efficient Transformer Training via Hybrid Normalization

### v3

@@ Abstract
Transformers have become the de facto architecture for a wide range of machine learning tasks, particularly in large language models (LLMs). Despite their remarkable performance, challenges remain in training deep transformer networks, especially regarding the position of layer normalization. While Pre-Norm structures facilitate more stable training owing to their stronger identity path, they often lead to suboptimal performance compared to Post-Norm. In this paper, we propose HybridNorm, a simple yet effective hybrid normalization strategy that integrates the advantages of both Pre-Norm and Post-Norm. Specifically, HybridNorm employs QKV normalization within the attention mechanism and Post-Norm in the feed-forward network (FFN) of each transformer block. We provide both theoretical insights and empirical evidence demonstrating that HybridNorm improves gradient flow and model robustness. Extensive experiments on large-scale transformer models, including both dense and sparse variants, show that HybridNorm consistently outperforms both Pre-Norm and Post-Norm approaches across multiple benchmarks. These findings highlight the potential of HybridNorm as a more stable and effective technique for improving the training and performance of deep transformer models.
@@ 1 Introduction
In transformers, Layer Normalization (LayerNorm) [1] plays a central role in stabilizing training by normalizing the activations within each layer. The two predominant strategies for applying LayerNorm are Pre-Layer Normalization (Pre-Norm) and Post-Layer Normalization (Post-Norm), each with its respective benefits and trade-offs. In the Pre-Norm architecture, normalization is applied before the residual addition, resulting in a more prominent identity path that facilitates faster convergence and more stable gradients [43]. This design is particularly advantageous when training deep models, as it helps mitigate gradient-related issues that can arise during backpropagation. However, while Pre-Norm can stabilize training, it often leads to inferior final performance compared to Post-Norm [42]. In contrast, Post-Norm applies normalization after the residual connection, resulting in stronger regularization effects, which contribute to improved model performance. This approach has been shown to improve the generalization ability of transformers, particularly in very deep networks [39]. Further discussion of related work is provided in Appendix 7..
@@ 2 Preliminaries
where $\mathrm{head}_{i}=\mathrm{attn}(Q_{i},K_{i},V_{i})$ for $i=1,2,\dots,h$, $\{\bullet_{i}\}_{i=1}^{h}=\mathrm{Split}(XW_{\bullet})$ for $\bullet\in\{Q,K,V\}$, and learnable parameters $W_{Q},W_{K},W_{V},W_{O}\in\mathbb{R}^{d\times d}$ By enabling the model to focus on different subspaces of the input representation, MHA enhances the transformer’s capacity to capture diverse patterns in the input sequence.
@@ 4.1 Benefits of Hybrid Method
where the gradient of normalization is $\frac{\partial\mathrm{Norm}(x)}{\partial x}=\alpha\odot\frac{\sqrt{d}}{||x||_{2}}\left(I-\frac{xx^{\top}}{||x||_{2}^{2}}\right)$. The gradient of Post-Norm is the product of two gradients, one of which is the normalization. If the spectral radius of the normalization gradient is less than 1, it causes gradient vanishing. In Pre-Norm, the residual connection is isolated from the normalization, ensuring that gradients retain a lower bound, thereby preventing gradient vanishing. To ensure relatively stable gradients, a natural idea is to use both types of normalization within a Transformer block, leading to Pre-Post (Pre-Norm in MHA and Post-Norm in FFN) and Post-Pre. From Table 4, we find that Pre-Post achieves the best performance, which is why we adopt this. And placing Pre-Norm in MHA with QKV-Norm further enhances performance.
@@ 4.2 Benefits of QKV-Norm
The above theorem is an informal version of Theorem 2, with a more precise statement and proof provided in Appendix 8. In attention with Pre-Norm, gradients of weights exhibit strong dependencies on other weights; for instance, $W_{Q}$ and $W_{K}$ are influenced by all three other weights but not by themselves. In contrast, with QKV-Norm, the gradient of each weight depends at most on itself and $W_{O}$. This indicates that the gradient in Pre-Norm is more tightly coupled with other weights compared to QKV-Norm, while Pre-Norm with QK-Norm lies between the two. Consequently, during training, if the norm of a certain weight becomes excessively large, it is harder to control in Pre-Norm, leading to increased gradient magnitude. This creates a vicious cycle that may cause model collapse. In contrast, QKV-Norm alleviates this issue and significantly improves training stability. In summary, the degree of gradient coupling follows: Pre-Norm $>$ Pre-Norm with QK-Norm $>$ QKV-Norm; whereas training stability follows the reverse: Pre-Norm $<$ Pre-Norm with QK-Norm $<$ QKV-Norm.
@@ 5.1 Experiment Settings
Methods BasicArithmetic HellaSwag SciQ ARC-C ARC-E PIQA OQA COPA Avg.$\uparrow$
@@ 5.2 Main Results
We evaluate the performance of HybridNorm and HybridNorm on 1.2B dense transformer models. Figure 3 compares the training dynamics of dense models with different normalization methods. As shown in the figure, models with HybridNorm and HybridNorm exhibit consistently lower training loss and validation perplexity throughout training compared to Pre-Norm, highlighting their effectiveness in enhancing training stability and convergence. Table 1 presents the downstream evaluation results. HybridNorm consistently outperforms Pre-Norm in most tasks, achieving the highest average score. Notably, it demonstrates substantial improvements in tasks such as BasicArithmetic (+3.11), HellaSwag (+1.71), and COPA (+3.78), indicating enhanced generalization and robustness. These results underscore the scalability of HybridNorm in larger transformer models, further validating its effectiveness in improving both training stability and downstream performance. More results can be found in Figure 9. In Appendix 10.1, we present additional comparisons with other approaches, such as Post-Norm and Mix-LN. We further provide analyses of signal propagation and entropy dynamics across different methods, as detailed in Appendix 10.2 & 10.3.
@@
Methods HellaSwag ARC-C ARC-E PIQA WinoGrande OQA BoolQ COPA Avg.$\uparrow$
@@ 6 Conclusion
In this paper, we have presented HybridNorm, a novel hybrid normalization approach that effectively integrates the strengths of Pre-Norm and Post-Norm to address the longstanding trade-offs in transformer training. We have provided both theoretical and empirical insights into how HybridNorm stabilizes gradient flow while maintaining strong regularization effects, demonstrating its potential to improve both convergence and final model performance. Extensive experiments across diverse benchmarks validate the efficacy of our approach, with results consistently showing that HybridNorm outperforms traditional normalization strategies in terms of stability and accuracy. Our findings underscore the necessity of re-evaluating normalization position within transformer architectures, thereby laying the groundwork for further exploration of hybrid strategies. We posit that HybridNorm signifies a substantial advancement in the design of more robust and performant deep transformer models, offering practical advantages for the training of next-generation large-scale neural networks.

### v4

@@ Abstract
Transformers have become the de facto architecture for a wide range of machine learning tasks, particularly in large language models (LLMs). Despite their remarkable performance, many challenges remain in training deep transformer networks, especially regarding the position of the layer normalization. While Pre-Norm structures facilitate more stable training owing to their stronger identity path, they often lead to suboptimal performance compared to Post-Norm. In this paper, we propose HybridNorm, a simple yet effective hybrid normalization strategy that integrates the advantages of both Pre-Norm and Post-Norm. Specifically, HybridNorm employs QKV normalization within the attention mechanism and Post-Norm in the feed-forward network (FFN) of each transformer block. We provide both theoretical insights and empirical evidence to demonstrate that HybridNorm improves the gradient flow and the model robustness. Extensive experiments on large-scale transformer models, including both dense and sparse variants, show that HybridNorm consistently outperforms both Pre-Norm and Post-Norm approaches across multiple benchmarks. These findings highlight the potential of HybridNorm as a more stable and effective technique for improving the training and performance of deep transformer models.
@@ 1 Introduction
In transformers, Layer Normalization (LayerNorm) [1] plays a central role in stabilizing training by normalizing the activations within each layer. The two predominant strategies for applying LayerNorm are Pre-Layer Normalization (Pre-Norm) and Post-Layer Normalization (Post-Norm), each with its respective benefits and trade-offs. In the Pre-Norm architecture, normalization is applied before the residual addition, resulting in a more prominent identity path that facilitates faster convergence and more stable gradients [44]. This design is particularly advantageous when training deep models, as it helps mitigate gradient-related issues that can arise during backpropagation. However, while Pre-Norm can stabilize training, it often leads to inferior final performance compared to Post-Norm [43]. In contrast, Post-Norm applies normalization after the residual connection, resulting in stronger regularization effects, which contribute to improved model performance. This approach has been shown to improve the generalization ability of transformers, particularly in very deep networks [39]. Further discussion of related work is provided in Appendix 7.
@@ 2 Preliminaries
[…]
@@ 4.1 Benefits of Hybrid Method
[…]
@@ 4.2 Benefits of QKV-Norm
The above theorem is an informal version of Theorem 2, with a more precise statement and proof provided in Appendix 8. In attention with Pre-Norm, gradients of weights exhibit strong dependencies on other weights; for instance, $W_{Q}$ and $W_{K}$ are influenced by all three other weights but not by themselves. In contrast, with QKV-Norm, the gradient of each weight depends at most on itself and $W_{O}$. This indicates that the gradient in Pre-Norm is more tightly coupled with other weights compared to QKV-Norm, while Pre-Norm with QK-Norm lies between the two. Consequently, during training, if the norm of a certain weight becomes excessively large, it is harder to control in Pre-Norm, leading to increased gradient magnitude. This creates a vicious cycle that may cause model collapse. In contrast, QKV-Norm alleviates this issue and significantly improves training stability. In summary, the degree of gradient coupling follows: Pre-Norm >> Pre-Norm with QK-Norm >> QKV-Norm; whereas training stability follows the reverse: Pre-Norm << Pre-Norm with QK-Norm << QKV-Norm.
@@ 5.1 Experiment Settings
[…]
@@ 5.2 Main Results
We evaluate the performance of HybridNorm and HybridNorm on 1.2B dense transformer models. Figure 3 compares the training dynamics of dense models with different normalization methods. As shown in the figure, models with HybridNorm and HybridNorm exhibit consistently lower training loss and validation perplexity throughout training compared to Pre-Norm, highlighting their effectiveness in enhancing training stability and convergence. Table 1 presents the downstream evaluation results. HybridNorm consistently outperforms Pre-Norm in most tasks, achieving the highest average score. Notably, it demonstrates substantial improvements in tasks such as BasicArithmetic (+3.11), HellaSwag (+1.71), and COPA (+3.78), indicating enhanced generalization and robustness. These results underscore the scalability of HybridNorm in larger transformer models, further validating its effectiveness in improving both training stability and downstream performance. More results can be found in Figure 9. In Appendix 11.1, we present additional comparisons with other approaches, such as Post-Norm and Mix-LN. We further provide analyses of signal propagation and entropy dynamics across different methods, as detailed in Appendix 11.2 & 11.3.
@@
5.3 More Experiment in 7B Dense Model
The experimental setup primarily follows the 7B model configuration outlined in [27] and the number of training tokens is 150B. To further enhance model stability, we introduce an additional normalization layer to the output of the attention module in 7B model, with its weight initialized to $1/\sqrt{2L}$ [41], where $L$ denotes the number of model layers. Table 3 presents the training loss and validation perplexity (PPL), while Table 4 reports the downstream evaluation metrics.
The experimental results demonstrate the clear superiority of HybridNorm over the traditional Pre-Norm approach in the 7B model. Firstly, HybridNorm achieves a lower training loss (2.430 vs. 2.469), indicating more efficient optimization during training. This improvement in training dynamics translates into consistently better performance across a range of language modeling benchmarks. Specifically, HybridNorm yields lower perplexity scores on all evaluated datasets, including C4, Books, Common Crawl, Wiki, and Wikitext 103. For instance, perplexity on the C4 dataset drops from 15.32 to 14.83, suggesting stronger generalization across both structured and unstructured corpora.
Moreover, HybridNorm shows consistent improvements on all downstream tasks, covering various domains such as arithmetic reasoning, commonsense QA, and natural language inference. Notably, performance on Basic Arithmetic improves significantly from 43.50% to 50.67%. Across the full suite of tasks, including ARC, PIQA, COPA, BoolQ, and WinoGrande, HybridNorm outperforms Pre-Norm in every case, leading to an overall increase in average accuracy from 60.61% to 63.06%.
@@ 5.3 More Experiment in 7B Dense Model
[…]
@@
Table 3: Training loss and perplexity (PPL) comparison of Pre-Norm and HybridNorm for 7B dense mdoel across multiple datasets. $C$ means Common Crawl.
Methods Loss C4 Books $C$ peS2o Reddit Stack Wiki Pile Wikitext
Pre-Norm 2.469 15.32 13.37 17.10 8.07 20.31 3.57 9.96 7.85 10.09
HybridNorm 2.430 14.83 12.77 16.77 7.67 19.65 3.40 9.34 7.81 9.16
Table 4: Downstream evaluation results of 7B dense models trained with Pre-Norm and HybridNorm. All numbers denote task accuracies (%). BA and OAQ mean Basic Arithmetic and Openbook QA, respectively.
Methods BA
Hella
Swag
SciQ ARC-C ARC-E PIQA OQA COPA
Wino
Grande
[…]
Pre-Norm 43.50 69.03 46.57 41.47 74.95 76.71 39.40 84.00 63.00 67.43 60.61
HybridNorm 50.67 70.77 47.44 43.82 75.82 78.93 43.77 86.01 63.32 70.06 63.06
@@ 6 Conclusion
In this paper, we have introduced HybridNorm, a novel hybrid normalization strategy that has seamlessly integrated the advantages of both Pre-Norm and Post-Norm, thereby addressing the longstanding trade-offs in transformer training. We have provided both comprehensive theoretical and empirical analyses to demonstrate how HybridNorm has stabilized gradient propagation while preserving strong regularization effects, ultimately improving both convergence speed and final model performance. Extensive experiments across diverse benchmarks have substantiated the effectiveness of our approach, consistently showing that HybridNorm has outperformed conventional normalization schemes in terms of stability and accuracy. These findings have highlighted the importance of re-examining the role and placement of normalization within transformer architectures, paving the way for further exploration of hybrid normalization paradigms. We believe that HybridNorm has marked a significant step forward in the development of more robust and efficient transformer models, offering practical advantages for training next-generation large-scale neural networks.